A proton of mass mp = 1.67 x 10-27kg and charge +e = 1.60 x 10-19 C isaccelerated from rest through an electric potential of 210 kV. Whatis its final velocity?
Answer #1
A proton of mass mp = 1.67 x 10-27kg and charge +e = 1.60 x 10-19 C isaccelerated from rest through an electric potential of 210 kV. Whatis its final velocity?
energy the proton acquires=1/2mv*v=210000* 1.60 x10-19
so velocity=√2*210000* 1.60 x 10-19/1.67 x10-27=6.3434*106m/sec